OpenStudy (anonymous):

MEDALLLL PLZZ HELP! SOLVE FOR X: 3/4(x-12)=12

OpenStudy (p0sitr0n):

you need to isolate x's from constant terms (this is how we solve equations) Thus, divide by 3/4 => x-12=16 x=28

OpenStudy (anonymous):

So the answer is 28 @P0sitr0n

OpenStudy (anonymous):

Can you help me with 1 more problem? @P0sitr0n

OpenStudy (anonymous):

What do you need help with?

OpenStudy (anonymous):

3/4y-12=12 solve for y? @taylorz115820

OpenStudy (anonymous):

Is it the same thing as for x?

OpenStudy (anonymous):

Isolate y

OpenStudy (anonymous):

do you know what to do 1st?

OpenStudy (anonymous):

divide?

OpenStudy (anonymous):

No you isolate the variable last after you have everything else away from it Ex. 2x - 4 = 4 To love this equation you would first +4 to both sides in order to get 2x=8 then you would divide 2 from both sides to get the answer of x=4

OpenStudy (anonymous):

Make sense?

OpenStudy (anonymous):

yea but I always get confused whenever there's a fraction? @taylorz115820

OpenStudy (anonymous):

A fraction where?

OpenStudy (anonymous):

x = 28

OpenStudy (anonymous):

We are on a different problem now

OpenStudy (anonymous):

oh what is it

OpenStudy (anonymous):

3/4y-12=12 solve for y

OpenStudy (anonymous):

Trying to get her to understand when to add or multiply or whatever 1st

OpenStudy (anonymous):

y in (-oo:+oo) (3/4)*y = 12 // - 12 (3/4)*y-12 = 0 3/4*y-12 = 0 // + 12 3/4*y = 12 // : 3/4 y = 12/3/4 y = 16 y = 16

OpenStudy (anonymous):

I think you did it wrong why is there a 0 it is supposed to be 12

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

So whats the answer?

OpenStudy (anonymous):

(3/4)y-12=12 +12+12 (3/4)y = 24 Divide by the reciprocal of 3/4 which is 4/3 which then you would really be multiplying 4/3•24 which = 32

OpenStudy (anonymous):

So y = 32

OpenStudy (anonymous):

So the answer is 32?

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

and if you plug 32 for y you should get 12

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

okay thanks both of you guys :)